# Safe and Responsible Large Language Model : Can We Balance Bias Reduction and Language Understanding in Large Language Models?

### v3

@@ Abstract
Large Language Models (LLMs) have significantly advanced various NLP tasks. However, these models often risk generating unsafe text that perpetuates biases. Current approaches to produce unbiased outputs from LLMs can reduce biases but at the expense of knowledge retention. In this research, we address the question of whether producing safe (unbiased) outputs through LLMs can retain knowledge and language understanding. In response, we developed the Safety and Responsible Large Language Model (SR), an LLM that has been instruction fine-tuned on top of already safe LLMs (e.g., Llama2 or related) to diminish biases in generated text. To achieve our goals, we compiled a specialized dataset designed to train our model in identifying and correcting biased text. We conduct experiments, both on this custom data and out-of-distribution test sets, to show the bias reduction and knowledge retention. The results confirm that SR outperforms traditional fine-tuning and prompting methods in both reducing biases and preserving the integrity of language knowledge. The significance of our findings lies in demonstrating that instruction fine-tuning can provide a more robust solution for bias reduction in LLMs. We have made our code and data available at Safe-LLM.
@@
Bias, Large Language Models , Safe Language , Instruction Fine-tuning[inst1]organization=Vector Institute for Artificial Intelligence,city=Toronto, postcode=M5G 1M1, state=Ontario, country=Canada
[inst2]organization=University of Limerick,city=Castletroy, postcode=V94 T9PX, state=Limerick, country=Ireland
[inst2]organization=Sheridan College,city=Trafalgar Rd, postcode=L6H 2L1, state=Ontario, country=Canada
1 Introduction
@@
Large Language Models (LLMs) have greatly advanced Natural Language Processing (NLP) in areas such as text generation, machine translation, question-answering systems, and more [1]. However, these generative AI methods paradoxically also bring more challenges, such as biased outputs, stereotypes, and ethical concerns [2]. For example, a scenario where a neutral prompt is provided to LLMs such as ChatGPT and Copilot:
The generated output was as follows:
@@
This output demonstrates a clear gender bias by assuming that a successful CEO is typically a man, which reinforces gender stereotypes. For the sake of this study, we define “bias" as discrimination, toxicity, stereotypes, and harm in linguistic outputs or general language use, as discussed in related literature [3].
@@
Recent research into LLM evaluations [4, 5, 6, 2, 7] show that LLMs can unintentionally echo negative stereotypes related to gender, race, and other demographics [8, 9]. Such biases in LLM generations also tend to amplify with the expansion of model sizes [10, 11]. To address biased outputs from LLMs, initial strategies have focused on implementing guardrails to monitor and adjust user interactions with LLM applications [12]. Additional techniques include Red-teaming, which uses simulated attacks to evaluate and strengthen system robustness against bias [8]. During the fine-tuning stage, advanced methods such as Reinforcement Learning from Human Feedback (RLHF) and context distillation are used to refine model responses [13, 14, 15]. Additionally, some approaches incorporate adversarial demonstrations to prepare models against potential malicious attacks [16, 10].
@@
Some works prompt LLMs to critically examine their own biases [17], while others employ techniques such as output filtering, ranking, and calibration [18]. Techniques like data augmentation and balancing, as well as embedding-based, probability-based, and generated text-based debiasing, are also used [19]. The primary goal of these approaches is to encourage self-reflection, identify and amend biased content, and adjust probabilities to reduce bias in LLM outputs.
Despite these efforts, completely eliminating bias is a complex challenge, and intensive mitigation strategies can lead to overfitting, risking the loss of language understanding or knowledge [20]. While prompt engineering is resource-efficient, research indicates that fine-tuning LLMs often yields better results [20]. Thus, integrating both strategies—prompt engineering and fine-tuning—might offer a more comprehensive approach to effectively reduce biases in LLM outputs.
State-of-the-art LLMs like Llama2 [21] and Mistral-v0.2 [22] are already fine-tuned for safety through demonstrations and RLHF. Adding an extra layer of instruction fine-tuning on top of these models could further enhance their ability to handle nuanced tasks such as unbiased content generation and ethical decision-making [3]. In this work, we aim to explore the effectiveness of this additional fine-tuning layer in our work.
@@
1.
2.
RQ2: Does instruction-based fine-tuning on top of already safe models enhance their ability to handle custom tasks (e.g., bias mitigation) without compromising their integrity?
@@
To create a LLM capable of processing biased or unsafe111Throughout this paper, we use the term ”unsafe” for texts that are biased, toxic, harmful, carry stereotypes, or convey negative sentiments. Conversely, we refer to benign or debiased texts as ”safe.” content and generating an unbiased or safe version of it.
@@
To ensure that the process of generation safe variations of content does not compromise the natural language understanding capability of our system.
@@
To implement instruction-based fine-tuning on top of already safe models (e.g., Llama2) to enhance their adaptability and effectiveness for our custom task.
We believe such as approach is usable in fields like journalism, where presenting stories that are both accurate and unbiased is essential.
The primary focus of this study is the development and implementation of methods to produce safe variations of unsafe content, which is a language generation tasks rather than classification task. The primary contributions of our research are outlined as follows:
@@
We present a curated dataset of social media content containing potentially biased (unsafe) texts, along with unbiased (safe or benign) variations (counterparts) that are prepared by our team of subject matter experts.
@@
We introduce Safety and Responsible Large Language Model (SR), a instruction fine-tuned LLM built on top of the already safe Llama2-7B-Chat model [23], to instruction fine-tune the LLM on our custom dataset. This approach is generalizable to other LLMs, as we have made the code and data available for reproducibility and further research.
Empirical Analysis: The empirical analysis on multiple datasets, both within the training set, and out of distribution datasets (Toxigen, BOLD, StereoSet) demonstrate that our instruction fine-tuned model significantly reduces unsafe content more effectively than both smaller models as well as base LLMs with just prompts. The retention of knowledge in the LLM is confirmed through a focused experiment on language understanding and human evaluation.
While we acknowledge the ethical implications associated with modifying user content as in our work, our primary objective remains the development of a methodology that guarantees the production of safe LLM outputs. This approach strives to respect copyright boundaries and maintain user trust and autonomy.
@@ 2 Methodology
Our study focuses on generating safe variations of potentially unsafe content, which is language generation rather than content classification. Our SR framework is shown in Figure 1, which has a dataset layer and an instruction fine-tuning method.
@@
In this research, we define ‘bias’ as any content in generative AI that shows hate, toxicity, offensiveness, or discrimination, which might perpetuate stereotypes or unfair portrayals of specific groups based on age, gender, race, or religion [10, 24, 25]. The major risks associated with LLM outputs that we explore are: Bias, where LLMs may generate content favoring or disfavoring groups unfairly; Toxicity, which includes aggressive or offensive content such as hate speech, insults, and threats, compromising the respectfulness of online interactions [6]; Stereotyping, where LLMs propagate generalized, often inaccurate assumptions about groups or individuals, leading to non-diverse representations [26]; and Harm, where there is a risk of LLMs producing content that could incite violence or societal harm, undermining public safety and well-being [27].
2.1 Dataset Preparation
@@
For this study, we used a dataset from our extensive collection of approximately 3.7 million records222https://huggingface.co/datasets/newsmediabias/news-bias-full-data. This dataset includes diverse content from both news and social media platforms, all in English. It covers over 200 different bias aspects, such as ageism, sexism, and gender discrimination. From this collection, we selected a sample of 20,000 records, each with more than 100 words and featuring a mix of various bias aspects from our main dataset. We name this dataset as Content Moderation Dataset (CMD).
@@
Annotation Procedure: The annotation task involves evaluating each piece of text to identify whether it is a bias, toxicity, negative sentiment (stemming from stereotyping), and harm and label it. Additionally, to read each unsafe text and modify the texts to create benign or safe version. We assembled a diverse team of 15 volunteer annotators, including five experts from various fields (computer science, social science, psychology, and statistics), each mentoring three students. This team brings together a wide range of demographics and expertise. Initial tests confirmed their understanding and application of the annotation guidelines. Detailed annotation guidelines are provided in A.
@@
To determine the “gold" labels for LLM risks categories (bias, toxicity, negative sentiment, and harm) and for the safe variation, we initially used a majority vote. Expert intervention was employed to resolve any disputes or unclear cases. The consistency of the annotation process was evaluated using Fleiss’ Kappa [28], with scores ranging from 0.62 to 0.74 across different risk categories (as shown in A). This indicates substantial agreement. An average score of 0.78 was calculated for the benign variation, which also indicates a strong consensus and demonstrating the reliability of the annotations 333In our code, we have also provided an LLM-based annotation strategy for safe text generation to enhance scalability and ensure the reliability of outputs..
@@
Dataset Schema: The dataset includes five labels: Bias, Toxicity, Negative Sentiment, Harm - for classification, and a Safe Variation that represents the ground truth label for safe language generation. We made our dataset available on Huggingface 444https://huggingface.co/datasets/newsmediabias/instruction-safe-llm. The dataset schema is given below:
@@
The descriptive statistics of the CMD data are in B. Few annotated examples are given in Table 2 555Disclaimer: For the sake of examples, these are truncated versions of the original data..
@@
2.2 Model Architecture
At the core of our work is the Llama2-7B-chat model [23], chosen for its ability to generate safe language. The base Llama2-7B-chat model is already fine-tuned with safety features such as RLHF, context distillation, and safe demonstrations, details in original paper [23]. We added another layer of instruction fine-tuning to create SR by further fine-tuning it on our custom dataset. We chose the smallest variant, Llama2-chat-7B, for SR, prioritizing ease of use and cost-effectiveness for inference and deployment.
@@ 2.2 Model Architecture
Instruction Design We adapt our CMD to adhere to the Alpaca instruction dataset format (Stanford format for instruction fine-tuned datasets) [29]. This instruction format follows structure as:
# <s> <<SYS>> System prompt <<</SYS>> [INST] User prompt [/INST] Model Response </s>
@@
Where, <<SYS>>: Represents the system prompt. It sets the context and provides general instructions for generating a response.
@@
[INST]: Denotes the instruction segment specific to each conversation turn. User prompts, which may contain both safe and unsafe examples, are accompanied by this prefix.
User prompt: The user’s input or query.
Model Response: The generated response by the model based on the given system prompt and user prompt.
An example for this is given below and the usage in Figure 2.
@@
<s><<SYS>> You are a helpful assistant. <<</SYS>> [INST] Respond to the user’s post by providing a version that maintains the core message but is presented in a benign and unbiased manner. Ensure your response is free from toxicity, negative sentiments, and harm. The goal is to reframe the content to retain its original intent while ensuring it contributes positively to the conversation: "Just saw the latest movie, and it was terrible! Waste of time" [/INST] I appreciate your feedback! Movie preferences can be subjective, and it is okay to have different opinions. If you like, I can recommend some other films you might enjoy. </s>
Efficient Fine-Tuning with QLoRA We utilized Quantized Language Model Optimization with Low Rank Adapters (QLoRA)[30], a Parameter-Efficient Finetuning (PEFT) technique via bitsandbytes and HuggingFace transformers Trainer, to finetune the Llama2-7B-Chat model on our custom instruction dataset in the creation of SR . QLoRA significantly minimizes the memory requirements for obtaining strong fine-tuning results, and strikes a balance between precision and resource efficiency using 4-bit NormalFloat (NF4) representation, double quantization, and paged optimizers. More details on QLoRA is in C.
3 Experiments
3.1 Training Details and Hyper-Parameters
The SR model is fine-tuned on a single A40 GPU with support from 4 CPU cores, employing PEFT and 4-bit quantization via QLORA (Rank=64, alpha=16, dropout=0.2) to manage GPU memory limits. Training was constrained to 1 epoch (with trials up to 5) to avoid over-fitting, similar to base Llama2-7B experiences. We used a batch size of 16 for training, 8 for evaluation, saved checkpoints every 25 steps with an early stopping after 3, and set the learning rate to 2e-4 and utilized paged AdamW optimizer “paged_adamw_32bit”[30]. The max sequence length was limited to 1024 for faster inference, as well as greedy decoding strategy . The detailed hyper-parameters are in Table 3 and Table 4.
| Section | Parameter | Value |
| General | lora_r | 64 |
| General | lora_alpha | 16 |
| General | lora_dropout | 0.2 |
| General | task_type | CAUSAL_LM |
| General | bias | None |
| Bits and Bytes | use_4bit | True |
| Bits and Bytes | bnb_4bit_dtype | float16 |
| Bits and Bytes | bnb_4bit_quant | nf4 |
| Bits and Bytes | use_nested_quant | True |
Other details as:
use_4bit = True % Activate 4-bit precision base model loading.
bnb_4bit_compute_dtype = "float16" % Compute dtype for 4-bit base models.
Note: bnb_4bit_compute_dtype for merging adapter+base model after finetuning.
bnb_4bit_quant_type = "nf4" % Quantization type (fp4 or nf4).
use_nested_quant = True % Activate nested quantization for 4-bit base models (double quantization).
compute_dtype = getattr(torch, bnb_4bit_compute_dtype).
| Parameter | Value | Parameter | Value |
| num_epochs | 1 | adam_beta1 | 0.9 |
| fp16 | Yes | adam_beta2 | 0.999 |
| bf16 | No | adam_epsilon | - |
| batch_size | 16/8 | training steps | 25 |
| max_grad_norm | 0.3 | grad_accum_steps | 1 |
| lr | 2e-4 | compute | 1xA40, 4xCPUs |
| optimizer | paged_adamw | memory | 100GB |
| scheduler | constant | runtime | 50m |
| warmup_ratio | 0.03 | ||
| weight_decay | 0.001 | ||
| seq_length | 1024 |
Carbon Footprint : To measure the environmental impact of training the SR model, the PEFT setup using one A40 GPU and four CPUs for 50 minutes had an energy use of 0.53 kWh and emitted 0.21 kgCO2e . This carbon footprint [31] is notably low, especially when contrasted with more demanding tasks, such as a dense (full) fine-tuning, or training Llama2, which produced 539 tCO2eq, fully offset by Meta’s sustainability efforts. The calculations for carbon footprinting are given in D.
3.2 Evaluation Datasets

### v5

@@ Abstract
Large Language Models (LLMs) have advanced various Natural Language Processing (NLP) tasks, such as text generation and translation, among others. However, these models often generate texts that can perpetuate biases. Existing approaches to mitigate these biases usually compromise knowledge retention. This study explores whether LLMs can produce safe, unbiased outputs without sacrificing knowledge or comprehension. We introduce the Safe and Responsible Large Language Model (SR), which has been instruction fine-tuned atop of a safe fine-tuned auto-regressive decoder-only LLM to reduce biases in generated texts. We developed a specialized dataset with examples of unsafe and corresponding safe variations to train SR to identify and correct biased text. Experiments on our specialized dataset and out-of-distribution test sets reveal that SR effectively reduces biases while preserving knowledge integrity. This performance surpasses that of traditional fine-tuning of smaller language models and base LLMs that merely reply on prompting techniques. Our findings demonstrate that instruction fine-tuning on custom datasets tailored for tasks such as debiasing is a highly effective strategy for minimizing bias in LLM while preserving their inherent knowledge and capabilities. The code and dataset are accessible at SR-LLM.
@@
keyword1, Keyword2, Keyword3, Keyword41 Introduction
@@
Large Language Models (LLMs) have greatly advanced Natural Language Processing (NLP) in areas such as text generation, machine translation, question-answering systems, and more [1]. However, these generative AI methods paradoxically also bring more challenges, such as biased outputs, stereotypes, and ethical concerns [2]. For example, this example demonstrates how a neutral prompt, when given to LLM such as ChatGPT and Copilot, can yield outputs that inadvertently reflect gender bias. Below, we present the prompt and the averaged response from multiple trials from these models.
The generated responses on average are depicted below:
@@
This output demonstrates a clear bias by assuming that a successful CEO is typically a man, which reinforces gender stereotypes. In the context of this study, we define “bias" as the presence of unsafe text that includes discrimination, toxicity, stereotypes, or harm within linguistic outputs or broader language usage, as explored in the relevant literature. [3].
@@
Recent research into LLM evaluations [4, 5, 6, 2, 7] also show that LLMs can unintentionally echo negative stereotypes related to gender, race, and other demographics [8, 9]. To address biased outputs from LLMs, initial strategies have focused on implementing guardrails to monitor and adjust user interactions with LLM applications [10]. Additional techniques include Red-teaming, which uses simulated attacks to evaluate and strengthen system robustness against biases [8]. During the fine-tuning stage, advanced methods such as Reinforcement Learning from Human Feedback (RLHF) and context distillation are used to refine model responses [11, 12, 13]. Additionally, some approaches incorporate adversarial demonstrations to prepare models against potential malicious attacks [14, 15].
@@
Related works on mitigating biases in texts prompt LLMs to critically examine their own biases [16], while some methods are based on techniques such as output filtering, ranking, and calibration [17]. Techniques like data augmentation and balancing, as well as embedding-based, probability-based, and generated text-based debiasing, are also used [18]. Research shows that while prompt engineering is generally resource-efficient, fine-tuning LLMs for various tasks often yields better results [19]. The primary goal of all these approaches is to encourage self-reflection, identify and amend biased content, and adjust probabilities to reduce bias in LLM outputs.
Despite significant recent efforts to mitigate bias in language generation, the complete elimination of bias presents a complex and ongoing challenge. Intensive mitigation strategies can often lead to overfitting, which occurs when a model becomes too specialized to the training data and fails to generalize well to new, unseen data [20]. This overfitting risks the loss of language understanding or knowledge retention because the model may focus too much on reducing bias at the cost of retaining the broader context and nuances of language [19].
State-of-the-art LLMs such as Llama2/3 [21], the Mistral-series [22], Claude, Gemini, and others [1] are increasingly fine-tuned by their developers and contributors to incorporate safety guardrails. These safety mechanisms often include techniques such as RLHF, context distillation, and curated safe demonstrations. Recent research suggests that leveraging fewer task-specific demonstrations with these safety-optimized models can further mitigate biases in generated outputs [16, 23]. However, it is worth considering whether an additional layer of instruction-based fine-tuning, employing targeted prompts and demonstrations, could further enhance the ability of these models to handle nuanced tasks such as unbiased content generation [3]. Thus, integrating both strategies, prompt engineering and fine-tuning, may provide a more comprehensive approach to effectively reducing biases in LLM outputs. In this work, we aim to explore the effectiveness of this additional fine-tuning layer.
The dataset and approach used in this paper build upon our seminal previous works [24, 25], incorporating a larger dataset and refined methodologies. Key differences between the current study and our prior work include fine-tuning a diverse set of models (e.g., Llama series and other instruction-tuned models), conducting experiments on out-of-distribution datasets, and performing both qualitative and quantitative evaluations. The qualitative analysis involved human expert reviews, while the quantitative analysis included beyond-accuracy evaluations through fairness metrics such as language proficiency, language modeling scores, stereotypes, content diversity, style evaluation, and statistical validation. This version also explores dense fine-tuning alongside quantization, focusing on the impact of these methods on accuracy and efficiency. These enhancements ensure robustness and reliability in our proposed approach.
Research questions
RQ2: To what extent does fine-tuning outperform few-shot and zero-shot prompting in reducing specific types of bias and retaining task-specific knowledge in LLMs?
RQ3: Does instruction-based fine-tuning on top of inherently safe models enhance their ability to handle custom tasks (e.g., bias mitigation) without compromising their integrity?
@@
Develop a safe and responsible LLM capable of identifying biased or harmful content and transforming it into a safe, unbiased version 111 In this context, “Unsafe” refers to texts that are biased, toxic, harmful, carry stereotypes, or convey negative sentiments, while “safe” texts are benign or debiased..
@@
Ensure that the process of converting unsafe texts to safe versions (a language generation task) does not diminish the model natural language understanding capabilities.
@@
Apply targeted fine-tuning to autoregressive, decoder-only LLMs with built-in safety guardrails to enhance their adaptability and effectiveness in generating safe text while preserving knowledge retention.
Contributions
This study focuses on the development and implementation of a safe and responsible LLM capable of generating safe variations of potentially unsafe content, emphasizing language generation rather than classification tasks. The key contributions of this research are as follows:
@@
We present a curated dataset of social media content containing potentially unsafe (biased) texts alongside unbiased (safe or benign) variations. This dataset, annotated by a team of twenty reviewers, including subject matter experts and the corresponding team, is specifically designed for benign text generation tasks.
@@
We introduce the Safe and Responsible Large Language Model (SR), an instruction fine-tuned LLM built on the Llama-instruct [26] architecture, tailored to our custom dataset for safe text generation. Our approach is generalizable to other instruction LLMs, and we have made the code and data publicly available to facilitate reproducibility and further research.
3.
We employ parameter-efficient fine-tuning methods, such as QLoRA [27], to optimize resource usage while maintaining high computational performance during training.
4.
We demonstrate the applicability of SR through a case study focused on job postings, a high-stakes domain where biased language can negatively impact diversity and inclusivity. SR showcases its potential for real-world applications in safe content generation and debiasing tasks.
The novelty of this work lies in the development of a custom dataset specifically designed for debiasing, incorporating paired biased and unbiased content to enable targeted instruction fine-tuning. Additionally, we establish a systematic pipeline that integrates LLM instruction design, fine-tuning, and evaluation, offering a cohesive framework for safe and responsible AI deployment. This approach is adaptable to high-stakes applications by leveraging domain-specific custom datasets, demonstrating its potential for practical and impactful real-world use cases. Empirical Analysis Our empirical analysis, conducted on both our training set and out-of-distribution datasets such as Toxigen, BOLD, and StereoSet, demonstrates the better performance of our instruction fine-tuned model, SR, in reducing unsafe content and retaining knowledge in the language generation task. This approach outperforms both smaller encoder-decoder fine-tuned language models and vanilla LLMs operating under zero-shot and few-shot prompt settings. The retention of important knowledge within the LLM is confirmed through targeted experiment on language understanding and human evaluation.
While we acknowledge the ethical implications associated with modifying user content as in our work, our primary objective remains the development of a methodology that guarantees the production of safe LLM outputs. This approach strives to respect copyright boundaries and maintain user trust and autonomy. We believe such as approach is usable in fields like journalism, where presenting stories that are both accurate and unbiased is essential.
@@ 2 Methodology
Problem definition
The core objective of our study is to address the challenge of generating safe variations of potentially unsafe content. This task involves language generation processes rather than content classification, positioning our research within the domain of creating responsible and safe text generation.
Preliminaries
@@
In this research, we define ‘bias’ as content in generative AI that exhibits hate, toxicity, offensiveness, or discrimination, which might perpetuate stereotypes or unfair portrayals of specific groups based on age, gender, race, or religion [15, 28, 29]. The major risks that we identify in this work with LLM outputs are: Bias, where LLMs may generate content favoring or disfavoring certain demographic groups (based on age, gender, race, religion, social status, etc.) unfairly; Toxicity, which includes aggressive or offensive content such as hate speech, insults, and threats, compromising the respectfulness of online interactions [6]; Stereotyping, where LLMs propagate generalized, often inaccurate assumptions about groups or individuals, leading to non-diverse representations [30]; and Harm, where there is a risk of LLMs producing content that could incite violence or societal harm, undermining public safety and well-being [31].
We employ the following terms for training models: Fine-tuning, which refers to adjusting the weights of a pre-trained model through additional training on a specific dataset to enhance its performance on related tasks; Prompts with demonstrations or few-shot learning, which involves providing the model with input-output example pairs (referred to as N-shots) along with prompts to guide accurate understanding and response generation [32]; and Instruction fine-tuning, which aims to improve a model’s ability to follow explicit instructions and respond appropriately by training it on a set of such instructions [33].
Our SR framework, illustrated in Figure 1, consists of a dataset layer and an efficient instruction fine-tuning method. The detailed steps of the framework are explained next.
2.1 Data preparation
@@
For this study, we utilized a subset of our extensive dataset, which comprises approximately 3.7 million records 222News Media Bias Full Data. This dataset encompasses a diverse array of content sourced from both news platforms and social media, all of which is in English. It covers a wide range of over 200 different bias aspects, including ageism, sexism, and gender discrimination.
From this comprehensive dataset, we specifically selected a sample of 20,000 records to form the Content Moderation Dataset (CMD). Each record in this sample contains more than 100 (upto 500) words, to ensure substantive content for analysis. We employed stratified sampling to ensure that all relevant categories of bias from our larger dataset were adequately represented. The purpose is to maintain a representative subset of the original dataset.
@@
Annotation Procedure The annotation task involves evaluating each piece of text to identify whether it is a bias, toxicity, negative sentiment (stemming from stereotyping), and harm and label it. The most important annotation task is to read each unsafe text and modify the texts to create benign or safe version; this is the ground truth label that we use in our model training and for evaluation.
We assembled a diverse team of 20 volunteer annotators: five experts from disciplines (computer science, social science, psychology, and statistics), each mentoring three students (master’s and one doctoral students) . This team brings together a range of demographics and expertise. Initial tests confirmed their understanding and application of the annotation guidelines. Detailed annotation guidelines are provided in A.
@@
To determine the labels for LLM bias categories (bias, toxicity, negative sentiment, and harm) and for the safe variation (we say as “gold" labels in this work), we used a majority vote. Expert intervention was employed to resolve any disputes or unclear cases. The consistency of the annotation process was evaluated using Fleiss’ Kappa [34], with scores ranging from 0.62 to 0.74 across different risk categories (as shown in A), which indicates substantial agreement. An average score of 0.78 was calculated for the benign variation, which also indicates a strong consensus and demonstrating the reliability of the annotations 333In our code, we have also provided an LLM-based annotation strategy for safe text generation to enhance scalability and ensure the reliability of outputs..
@@
Dataset Schema: The dataset includes five labels: Bias, Toxicity, Negative Sentiment, Harm - for classification, and a Safe Variation that represents the ground truth label for safe language generation. We made our dataset available on Huggingface 444https://huggingface.co/datasets/newsmediabias/instruction-safe-llm. The dataset schema is given below and detail in Table 1:
@@
The descriptive statistics of the CMD data are in B. Few annotated examples are given in Table 2 555Disclaimer: For the sake of examples, these are truncated versions of the original data.. The process of building the CMD dataset took about three months, including data selection, preprocessing, annotation, and validation of labels.
@@
2.2 Model
At the core of our work is an autoregressive, decoder-only Large Language Model (LLM) designed to handle instruction-following tasks with precision and coherence. We utilize Llama (instruction model) [26] for its proven architectural robustness and suitability for fine-tuned instruction-following applications, though other instruction-following models could also be employed. The base model is already optimized with essential safety features, including RLHF, context distillation, and curated safe demonstrations [26].. Building on this foundation, we introduced an additional layer of instruction fine-tuning to create SR, leveraging our custom dataset to enhance task-specific performance. Our methodology is versatile and can be extended to other models within the Llama series or similar instruction-tuned frameworks [35] such as Mistral-Instruct, Phi-Instruct, Gemma, and Qwen. These models follow the instruction-tuned paradigm, combining a base LLM with additional fine-tuning and RLHF to boost task-specific capabilities.
Training objective
To fine-tune , we adopt an autoregressive training paradigm. Specifically, we seek to minimize the cross-entropy loss:
where represents the model parameters, is the training dataset, denotes the instruction (and optional context), and is the corresponding target sequence. By iteratively optimizing this loss, the model learns to generate coherent, contextually relevant responses to diverse instruction prompts.
Instruction design
@@ 2.2 Model
We adapt our CMD dataset to adhere to the Alpaca instruction dataset format (Stanford format for instruction fine-tuned datasets) [36]. This instruction format follows the structure:
# <s> <<SYS>> System prompt <</SYS>>
[INST] User prompt [/INST]
Model Response </s>
Where,
[leftmargin=*]
@@
<<SYS>>: Represents the system prompt. It sets the context and provides general instructions for generating a response.
@@
[INST]: Denotes the instruction segment for each conversation turn. User prompts, which may contain both safe and unsafe examples, are preceded by this prefix.
User prompt: The user’s query or message.
Model Response: The model-generated response based on the system prompt and user prompt.
An example snippet is shown below, and its usage is illustrated in Figure 2:
@@
<s><<SYS>> You are a helpful assistant. <</SYS>> [INST] Reframe the user’s post to: - Retain the original intent and factual accuracy. - Present the content in a benign, unbiased, and positive manner. - Avoid toxicity, negative sentiments, or harmful language. - Contribute positively to the conversation. Example Input: "Just saw the latest movie, and it was terrible! Waste of time." Example Output: "I appreciate your feedback! It’s okay to have different opinions about movies. If you’d like, I can recommend some other films you might enjoy." [/INST] </s>
2.2.1 Efficient parameterization for scalable fine-tuning
In this work, we used parameter-efficient methods to enhance the performance of LLMs while reducing computational demands. The objective was to make easy to retrain and production-ready. We primarily utilized Quantized Low-Rank Adaptation (QLoRA) [27], a technique that reduces the memory footprint and computational requirements of fine-tuning LLMs by combining low-rank adapters with 4-bit quantization. Formally, low-rank adaptation can be viewed as a decomposition of parameter updates:
where is the dimensionality of the original parameter matrix , and is the low-rank dimension, allowing us to store far fewer parameters. Combining such decompositions with 4-bit NormalFloat (NF4) quantization further cuts down on computational and memory costs. More details on the QLoRA implementation can be found in Sections 3.1. After fine-tuning, the adapter weights were merged with the base model to ensure stability and prepare the model for production deployment. The resulting merged model weights are publicly available666https://huggingface.co/newsmediabias/UnBIAS-LLama2-Debiaser-Chat for use.
As an alternative, we also experimented with prefix-tuning [37], a method that optimizes continuous task-specific vectors (prefixes) while keeping the pre-trained model parameters fixed. By training only the prefix tokens, this approach enabled quick task-specific adaptations without retraining the entire model. Both methods demonstrated state-of-the-art performance in reducing computational overhead and maintaining model accuracy [38]. Our goal is to use these parameter-efficient techniques for creating lightweight yet effective LLM extensions tailored for domain-specific and debiasing tasks.
Our fine-tuning framework balances neutrality and expressiveness by combining curated datasets and tailored instructions. Harmful content, such as “This group is always lazy", is neutralized to “Such generalizations are harmful", while benign opinions, like “I enjoy painting" remain unchanged. Implicit guardrails, including RLHF, prevent over-correction, ensuring factual responses, such as “Wasting water is harmful" remain unaffected. Evaluations in Section 4 demonstrate above 92% success in neutralizing harmful bias and above 85% preservation of knowledge in benign cases.
3 Experimental Setup
3.1 Training details and hyperparameters
Hardware and Runtime: In our experiments, the SR model was trained using two different parameter-efficient techniques: QLoRA [27] and prefix-tuning [37]. Both methods were implemented on a single NVIDIA A100 GPU with support from 4 CPU cores. The total memory usage was approximately 100GB, and the total runtime was around 50 minutes for QLoRA and 30 minutes for prefix-tuning.
We used a batch size of 16 for training and 8 for evaluation. Training was constrained to 1 epoch for QLoRA (with trials up to 5 epochs; more epochs led to overfitting, as noted in the Llama2 paper [26]) and 5 epochs for prefix-tuning. The maximum sequence length was set to 1024 for faster inference. Early stopping was applied with a patience of 3, and checkpoints were saved every 25 steps.
We set the gradient accumulation steps to 1 and applied a maximum gradient norm (‘max_grad norm‘) of 0.3 to prevent gradient explosion. The learning rate scheduler was set to constant, with a warmup ratio of 0.03. A weight decay of 0.001 was used for regularization. We enabled FP16 precision (‘fp16=True‘) to reduce memory consumption, and BF16 precision was disabled (‘bf16=False‘). We used the AdamW optimizer with betas set to (‘beta1=0.9‘, ‘beta2=0.999‘).
Quantization Settings for QLoRA: For QLoRA, we set the LoRA rank () to 64, to 16, used a dropout rate of 0.2, and applied 4-bit NF4 quantization with nested quantization enabled. The compute data type was set to float16. We also specified the task type as CAUSAL_LM and set the bias to None.
To activate 4-bit precision base model loading, we set use_4bit to True. The compute data type for 4-bit base models was specified as float16 by setting bnb_4bit compute_dtype to "float16". We selected the quantization type as NormalFloat (NF4) by setting bnb_4bit quant_type to "nf4", which offers better precision compared to FP4. Nested quantization was enabled (use_nested_quant = True) to achieve double quantization, further reducing memory usage. The compute data type was set using compute_dtype = getattr(torch, bnb_4bit compute_dtype). The parameter bnb_4bit_compute_dtype is particularly important when merging the adapter and base model after fine-tuning.
| Hyperparameter | Value | Hyperparameter | Value |
| General Training Parameters | |||
| Batch size (training) | 16 | Batch size (evaluation) | 8 |
| Number of epochs (QLoRA) | 1 (trials up to 5) | Number of epochs (Prefix Tuning) | 5 |
| Max sequence length | 1024 | Evaluation steps | Every 25 steps |
| Early stopping patience | 3 | Gradient accumulation steps | 1 |
| Max gradient norm | 0.3 | Scheduler | Constant |
| Warmup ratio | 0.03 | Weight decay | 0.001 |
| FP16 enabled | Yes | BF16 enabled | No |
| Optimizer and Learning Rate | |||
| Optimizer (QLoRA) | Paged AdamW 32bit | Learning rate (QLoRA) | |
| Optimizer (Prefix Tuning) | AdamW | Learning rate (Prefix Tuning) | |
| Adam Beta1 | 0.9 | Adam Beta2 | 0.999 |
| QLoRA-Specific Hyperparameters | |||
| LoRA rank () | 64 | LoRA | 16 |
| LoRA dropout | 0.2 | Quantization type | 4-bit NF4 |
| Use nested quantization | True | Compute data type | float16 |
| Task type | CAUSAL_LM | Bias | None |
| Prefix-Tuning Specific Hyperparameters | |||
| Prefix length | 30 | Number of prefixes per layer | 2 |
| Dropout rate | 0.1 | Initialization method | Random Normal |
| Number of epochs | 5 | Learning rate | |
Prefix-Tuning Settings: For prefix-tuning, we trained only the prefix parameters while keeping the base model weights fixed. We trained for 5 epochs, setting the prefix length to 30 and optimizing the prefix vectors using a learning rate of . The prefix vectors were initialized using a random normal distribution, aiding in better convergence during training. We used the AdamW optimizer, set the number of prefixes per layer to 2, and used a dropout rate of 0.1. The detailed hyperparameters are provided in Table 3.
Carbon Footprint: To assess the environmental impact of training the SR model, we evaluated both QLoRA and Prefix-Tuning methods. For QLoRA, the PEFT setup used one A100 GPU and four CPUs for 50 minutes, consuming 0.53 kWh of energy and emitting 0.21 kgCO2e. Prefix-Tuning was more efficient, requiring only 30 minutes under the same hardware conditions, leading to an energy consumption of 0.32 kWh and carbon emissions of 0.13 kgCO2e. The formulas to calculate these emissions is taken from related work [39]. These footprints are small compared to resource-intensive tasks like dense (full) fine-tuning or training Llama2, which emitted approximately 539 tCO2e, as fully offset by Meta’s sustainability efforts. Detailed calculations, including emission factors and regional energy assumptions, are provided in D.
3.2 Evaluation data